OpenStudy (anonymous):

find the sum of the following arithmetic series 7+11+15+19+.....+67

OpenStudy (anonymous):

you're summing 16 terms, so it's (7 + 67)*16/2 = 74*8 = 148*4 = 296*2 = 592

OpenStudy (anonymous):

thank u